What is the pH of a solution with [H^+] of 1.0 x 10^9M? What is its pOH? (Kw = [H^+][OH^-])


1
Expert's answer
2022-05-12T06:17:02-0400

"pH=-ln [H^+]=-ln(1.0 x 10^9)=-9"

pH+pOH=14

pOH=14+9=23